# Excitation Spectrum and Spin Hamiltonian of the Frustrated Quantum Ising Magnet PrBWO

### v1

@@ VI Acknowledgements
The effective point-charge refinement to magnetometry and INS data was carried out with four fit parameters: The three charges from inequivalent O ligands caging the Pr ions, as well as a mean field exchange constant meV. The predicted CEF spectrum consists of nine isolated singlets, with excitations at , 37.0, 42.6, 58.0, 74.8, 91.8, 101.1 and 105.1 meV. Due to the K gap to higher CEF excitations, the lowest two levels form a well isolated quasi-doublet ground state with a small splitting meV. Nearly all of the neutron scattering intensity is concentrated on this transition, with further modes suppressed by a factor and hence not visible in experiment above noise level. The CEF parameters obtained from the fitting procedure are presented in Table 5, while the wavefunctions of the lowest quasi-doublet are given in Table 4. Since those two states are singlets, we have and the transverse matrix elements are Ising-like, with . Despite the perfect axial anisotropy ( and ) at a single-ion level, upon averaging over all six ions in the unit cell we obtain and , in excellent agreement with experiment.
Atom |
Pr |
0.0854(3) |
0.7245(3) |
0.3460(7) |
0.27(5) |
B |
0 |
0 |
0 |
0.49(9) |
W |
0.3333 |
0.6667 |
0.8819(9) |
0.23(7) |
O(1) |
0.0459(2) |
0.8689(2) |
0.9940(6) |
0.65(6) |
O(2) |
0.1413(2) |
0.5192(2) |
0.1067(6) |
0.49(4) |
O(3) |
0.1931(2) |
0.4759(2) |
0.6724(6) |
0.98(8) |
Appendix B: Crystal Structure
@@
The atomic coordinates obtained from the low-temperature neutron powder diffraction experiment discussed in Sec. III.9 are provided in Table 6. These are deemed to be more suitable for point-charge calculations than other refinements available in the literature [ashtarNewFamilyDisorderFree2020, krutkoStructuresNonlinearHexagonal2006], as the positions of the light O ligands directly responsible for the CEF environment are extremely difficult to refine accurately through x-ray diffraction, especially in the presence of the heavy tungsten.

### v2

@@ Appendix A Effective Point-Charge Model
An effective point-charge fit to the combined INS and magnetometry data was carried out using a least-squares refinement with four fit parameters: The charges of the three inequivalent O2- ligands caging the Pr3+ ions, as well as the mean field exchange constant meV. The CEF parameters obtained from the fitting procedure are presented in Table 4, while the wavefunctions of the lowest quasi-doublet are given in Table 5. The optimized CEF spectrum consists of nine isolated singlets, with excitations at , 37.0, 42.6, 58.0, 74.8, 91.8, 101.1 and 105.1 meV. Due to the K gap to higher CEF excitations, the lowest two levels form a well isolated quasi-doublet ground state with a small splitting meV. Nearly all of the neutron scattering intensity is concentrated on this transition, with further modes suppressed by a factor and hence not visible in experiment above noise level. This is exacerbated at high temperatures, where the combination of lifetime broadening effects and strongly increased phonon population makes it increasingly difficult to identify CEF transitions.
| Atom | Occ. | Equiv. | |||
| Pr | 0.0854(3) | 0.7245(3) | 0.3460(7) | 1.00(1) | 0.27(5) |
| 11B | 0 | 0 | 0 | 0.32(1) | 0.49(9) |
| W | 0.3333 | 0.6667 | 0.8819(9) | 0.33(1) | 0.23(7) |
| O(1) | 0.0459(2) | 0.8689(2) | 0.9940(6) | 1.06(1) | 0.65(6) |
| O(2) | 0.1413(2) | 0.5192(2) | 0.1067(6) | 1.07(1) | 0.49(4) |
| O(3) | 0.1931(2) | 0.4759(2) | 0.6724(6) | 1.06(1) | 0.98(8) |
Appendix B Pseudospin Projection
Due to the large K gap to higher CEF excitations, only the quasi-doublet states defined in Eq. (III.3) are populated at low temperatures and the thermodynamic/magnetic properties are fully determined by these low-energy degrees of freedom. Therefore, it becomes useful to “integrate out” the irrelevant states and calculate the projection of the magnetic moments onto the quasi-doublet subspace spanned by , given as
Here, defines the ground state projector and the local pseudospin operators are constructed in the standard framework [rauFrustratedQuantumRareEarth2019]
from the two thermally accessible crystal field states. We note that our choice of basis conforms to the convention common to pyrochlore oxides [rauFrustratedQuantumRareEarth2019, martinDisorderQuantumSpin2017a]. In principle, an equivalent formulation in terms of the and states is possible and would correspond simply to swapping our notation of and throughout the text. With this, the total magnetic dipole moment operator can be written as
where is the longitudinal -factor and (as is the case for all non-Kramers (quasi-)doublets). Despite the perfect axial single-ion anisotropy, upon averaging over all six ions in the unit cell we obtain and , in excellent agreement with experiment. Only contributes to the magnetic dipole moment, represented as either “up” or “down”. On the other hand, the transverse pseudospin components show up in the projected crystal field Hamiltonian
acting as a transverse field on the otherwise Ising-like moments and mixing our basis into the quasi-doublet ground states given in Eq. (III.3). Since those two states are singlets, we have and only the transverse matrix elements are nonzero - meaning that the neutron scattering cross section is purely inelastic. All in all, we can establish a direct mapping between the low-energy physics of our full CEF Hamiltonian Eq. (10) and that of Ising-like pseudospin moments in a strong transverse field (TFIM) as in Eq. (4).
Appendix C Crystal Structure
@@
Rietveld refinements based on the low-temperature neutron powder diffraction experiment discussed in Sec. III.9 were carried out in the hexagonal space group. The refined lattice parameters are and , while the atomic coordinates, site occupancy and equivalent isotropic displacement parameters [truebloodAtomicDisplacementParameter1996] are provided in Table 6. The coordinates obtained in this way are deemed to be more suitable for point-charge calculations than other refinements available in the literature [ashtarNewFamilyDisorderFree2020, krutkoStructuresNonlinearHexagonal2006], as the positions of the light O2- ligands directly responsible for the CEF environment are extremely difficult to refine accurately through x-ray diffraction, especially in the presence of the heavy tungsten. All structural considerations and bond angles/distances discussed in Sec. II.1 are based on these parameters.